OpenStudy (anonymous):

Solve the equation by first clearing the fractions. \[\frac{ 1 }{ 2 }f - 5=1\]

OpenStudy (jdoe0001):

multiply both sides by " 2 "

OpenStudy (jdoe0001):

that'd clear out the fractions, then simplify

OpenStudy (anonymous):

will it be negative 8 ? or no

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{2}f-5=1\implies \cancel{ 2 }\cdot \cfrac{1}{\cancel{ 2 }}f-2\cdot 5=2\cdot 1\implies f-10=2\implies ?\)

OpenStudy (anonymous):

that would equal to negative 12 ? correct ?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw